# Multi-CATE: Multi-Accurate Conditional Average Treatment Effect Estimation Robust to Unknown Covariate Shifts

### v1

@@ Results
Train Shift CForest S-learner DR-learner T-learner
size degree OS wOS OS wOS OS Ridge Tree OS wOS Ridge Tree
0 -0.59 -0.66 -2.25 -2.27 -2.28 -2.32 -2.42 -4.59 -4.6 -3.73 -4.4
0.25 -2.94 -1.19 -4.51 -3.91 -3.96 -3.38 -3.88 -6.21 -5.58 -4.39 -5.64
0.5 -5.21 -1.94 -6.68 -5.78 -5.14 -5.16 -5.36 -7.94 -6.79 -6.71 -8.1
0.75 -7.23 -2.63 -8.54 -7.44 -6.94 -5.77 -6.43 -9.59 -7.92 -8.42 -9.96
500 1 -8.57 -4.95 -10.07 -8.73 -8.58 -6.96 -7.66 -11.09 -9.13 -9.26 -10.79
1.25 -9.07 -5.13 -10.71 -9.44 -6.83 -6.66 -7.65 -11.62 -9.41 -10.17 -11.92
1.5 -9.56 -6.54 -11.14 -9.84 -8.59 -6.72 -8.72 -12 -9.49 -10.49 -12.25
1.75 -9.71 -5.74 -11.54 -10.19 -8.4 -7.83 -8.57 -12.48 -9.76 -11.2 -12.58
2 -10.04 -7.48 -11.7 -10.16 -9.39 -7.1 -8.24 -12.63 -9.29 -10.75 -12.02
0 0.18 0.17 -2.05 -2.06 -1.62 -1.63 -1.48 -3.51 -3.52 -3.3 -3.46
0.25 -2.01 -0.22 -3.93 -3.37 -2.49 -2.87 -3.04 -4.83 -4.28 -4.3 -4.78
0.5 -3.88 -0.64 -5.86 -4.88 -3.36 -3.81 -3.81 -6.53 -5.42 -6.2 -7.06
0.75 -5.55 -1.76 -7.54 -6.37 -4.39 -3.9 -4.55 -8.12 -6.72 -7.29 -8.26
2000 1 -6.5 -1.8 -8.64 -7.27 -5.98 -5.41 -5.71 -9.15 -7.28 -8.18 -9.26
1.25 -7.18 -2.12 -9.45 -8.04 -6.11 -5.29 -5.73 -9.95 -7.68 -8.98 -10.53
1.5 -7.6 -2.23 -9.74 -8.27 -6.07 -6.02 -6.41 -10.26 -7.87 -9.3 -10.63
1.75 -7.8 -3.98 -10.05 -8.81 -6.2 -5.67 -6.62 -10.55 -8.44 -9.95 -11.77
2 -7.87 -2.7 -10.19 -8.6 -6.84 -5.42 -6.29 -10.63 -7.7 -9.79 -11.07
0 0.13 0.16 -2.01 -2 -1.41 -1.46 -1.54 -3.17 -3.17 -3.02 -3.15
0.25 -1.64 -0.01 -3.64 -3.09 -2.29 -2.4 -2.35 -4.37 -3.86 -4.1 -4.39
0.5 -3.5 -0.4 -5.53 -4.53 -3.6 -3.47 -3.77 -6.08 -5.04 -5.9 -6.55
0.75 -4.77 -0.37 -6.87 -5.61 -4.15 -4.01 -4.16 -7.27 -5.73 -6.72 -7.8
3500 1 -5.83 -0.64 -8.12 -6.73 -5.24 -4.64 -4.55 -8.57 -6.61 -7.76 -8.58
1.25 -6.39 -0.74 -8.74 -7.25 -4.73 -5.2 -6.17 -9.08 -7.02 -8.4 -9.62
1.5 -6.7 -0.03 -9.06 -7.42 -6.11 -5.35 -5.37 -9.5 -7.08 -8.98 -10
1.75 -7 -0.6 -9.29 -7.63 -6.05 -5.85 -5.81 -9.68 -7.09 -8.8 -10.08
2 -7.17 -1.68 -9.58 -7.77 -6.13 -6.14 -6.2 -10.07 -7.46 -9.1 -9.93
0 0.21 0.22 -1.91 -1.89 -1.2 -1.35 -1.23 -3 -2.99 -2.96 -3.28
0.25 -1.47 0.09 -3.49 -2.94 -2.01 -2.14 -2.18 -4.14 -3.67 -3.93 -4.63
0.5 -3.05 0.05 -5.08 -4.04 -2.9 -2.94 -3.08 -5.53 -4.47 -5.01 -5.58
0.75 -4.53 -0.21 -6.68 -5.4 -3.9 -3.92 -4 -7.11 -5.58 -6.81 -7.51
5000 1 -5.38 0.18 -7.65 -6.25 -4.79 -4.55 -4.7 -8.02 -6.14 -7.51 -8.49
1.25 -5.99 -0.97 -8.28 -6.76 -4.67 -5.11 -4.95 -8.64 -6.45 -8.24 -9.2
1.5 -6.32 -0.21 -8.76 -6.99 -5.27 -4.71 -5.39 -9.11 -6.76 -8.62 -9.71
1.75 -6.42 -0.64 -8.83 -7.1 -5.64 -4.39 -5.24 -9.28 -6.77 -8.57 -9.49
2 -6.88 -2.61 -9.24 -7.47 -6.14 -5.33 -5.35 -9.68 -7.28 -9.12 -10.24
@@
Train Shift CForest S-learner DR-learner T-learner
size degree OS wOS OS wOS OS Ridge Tree OS wOS Ridge Tree
0 48.96 48.83 61.44 61.47 49.01 36.24 41.68 50.37 50.47 37.92 48.71
0.25 58.03 52.83 79.93 79.13 64.22 47.74 53.11 72.51 64.08 52.5 67.26
0.5 75.47 59.45 107.39 107.74 76.5 65.54 69.39 103.71 86.43 80.42 111.06
0.75 98.37 78.02 134.31 132.05 109.67 83.42 87.06 132.66 105.15 107.44 147.33
500 1 116.47 94.48 160.77 154.28 126.34 110.72 111.87 164.12 127.02 121.08 165.23
1.25 121.81 132.84 169.13 165.05 142.91 108.59 128.54 173.35 133.51 138.64 186.45
1.5 128.21 144.18 177.03 169.83 133.75 115.82 119.92 179.79 137.04 142.15 194.81
1.75 129.24 139.23 183.05 177.42 153.78 109.25 123.72 189.41 140.58 158.14 198.73
2 136.23 156.07 187.09 176.18 143.81 114.75 133.21 193.61 136.73 148.47 192.43
0 31.25 31.29 41.25 41.46 26.21 23.87 25.76 32.69 32.76 28.04 33.36
0.25 35.78 33.1 56.21 54.81 37.36 31.1 35.15 49.45 42.36 40.36 51.18
0.5 47.93 36.68 78.99 74.7 53.38 47.29 58.23 74.61 57.9 68 86.61
0.75 62.81 43.77 102.92 98.62 67.95 55.98 59.31 101.23 78.57 83.52 107.64
2000 1 71.52 50.4 118.4 112.52 76.93 65.76 73.73 117.39 86.09 96.77 125.88
1.25 78.65 65.43 131.65 125.92 75.88 68 82.42 130.8 91.11 111.73 148.89
1.5 83.4 87.66 134.23 128.7 80.8 77.14 94 134.91 95.45 113.56 149.28
1.75 84.4 80.99 139.06 138.6 86.48 81.35 96.41 139.34 104.81 125.67 171.48
2 85.39 117.33 139.58 135.01 85.56 82.29 84.27 139.53 90.69 120.89 155.25
0 26.41 26.48 34.62 34.77 22.3 20.45 22.49 27.94 27.94 24.09 28.1
0.25 29.68 27.89 48.19 46.15 29.57 27.95 28.28 42.62 36.47 36.65 43.35
0.5 40.64 32.07 69.51 64.46 40.27 44.18 42.4 66.82 52.49 62.16 76.77
0.75 50.62 34.89 86.45 80.05 55.44 50.07 52.88 83.88 60.45 72.47 96.5
3500 1 60.8 40.59 105.13 99.35 58.89 55.2 65.03 104.92 73.67 88.04 109.7
1.25 65.42 53.8 113.57 105.53 73.34 60.93 71.11 112.44 79.46 98.13 128.19
1.5 67.51 58.93 115.71 105.15 66.37 63.82 62.2 117.23 78.11 106 131.51
1.75 71.14 84.67 119.63 110.99 68.86 60.95 72.15 119.74 79.36 100.75 133.81
2 72.57 94.48 124.32 112.8 70.37 68.83 68.96 127.26 84.85 106.19 128.92
0 23.76 23.84 30.99 31.12 19.81 18.33 19.01 25.41 25.37 22.87 28.73
0.25 26.53 25.12 43.11 40.92 25.98 25.28 26.89 38.97 33.2 36.33 45.83
0.5 35.34 28.45 60.29 54.75 32.56 31.91 35.93 57.47 42.99 48.56 60.4
0.75 46.53 31.58 81.14 73.84 48.41 41.51 53.5 81.54 58.2 75.15 91.08
5000 1 53.35 35.06 94.19 86.83 53.36 52.64 59.22 94.05 64.56 83.78 105.81
1.25 58.87 41.09 102.34 93.53 54.68 54.36 56.8 102.96 68.77 94.5 117.64
1.5 61.59 51.86 109.6 96.23 57.74 59.66 65.16 110.6 73.29 99.92 125.89
1.75 61.85 64.77 108.83 97.83 65.31 62.42 68.62 111.54 71.68 97.3 120.56
2 67.63 65.96 116.97 105 67.87 60.4 78.9 119.41 81.65 107.55 134.97
@@
Train Shift CForest S-learner DR-learner T-learner CForest S-learner DR-l T-learner
size degree OS OS OS Ridge Tree OS Ridge Tree RCT wRCT RCT wRCT RCT RCT wRCT
0 -3.62 -4.41 -5.87 -3.91 -4.45 -7.43 -4.06 -4.23 0.01 0 0.83 0.82 -0.07 0.05 0.05
0.25 -3.75 -4.43 -5.1 -4.29 -3.88 -7.52 -4.18 -4.16 0.85 -0.64 1.54 1.25 0.64 0.89 0.41
0.5 -3.59 -4.39 -4.91 -4.73 -4.16 -7.32 -4.36 -3.89 1.87 -1.05 2.08 1.57 1.75 1.67 0.84
0.75 -3.41 -3.99 -4.4 -2.55 -3.6 -6.73 -4.24 -3.32 3.2 -1.37 2.7 1.88 2.56 2.67 1.17
500 1 -3.61 -4.38 -6.05 -2.71 -1.57 -7.53 -5.08 -3.25 5.01 1.15 3.5 2.13 1.71 4.13 2.17
1.25 -3.44 -4.12 -4.73 -3.54 -3.64 -7.13 -4.95 -2.56 6.56 0.48 4.43 2.39 3.17 5.47 3.83
1.5 -3.49 -4.16 -4.96 -3.61 -3.02 -7.31 -4.79 -1.77 7.89 2.07 5.16 2.53 4.42 6.26 3.97
1.75 -3.45 -4.18 -6.22 -3.01 -3.7 -6.98 -5.08 -1.35 8.81 5.85 6.01 2.65 5.95 7.47 5.69
2 -3.58 -4.36 -4.98 -3.1 -3.43 -7.42 -4.7 -0.95 9.5 2.38 6.6 2.83 7.1 8.21 6.07
0 -2.73 -4.23 -3.45 -1.89 -1.27 -5.95 -3.73 -3.61 0.01 0.12 0.78 0.81 0.08 0.03 0.06
0.25 -2.59 -4.07 -3.36 -1.23 -1.4 -5.78 -3.19 -3.17 0.79 -1.29 1.44 1.08 0.46 0.72 0.16
0.5 -2.73 -4.39 -3.53 -1.59 -1.65 -6.07 -3.76 -3.24 1.8 -2.16 2.07 1.57 1.25 1.55 0.6
0.75 -2.81 -4.4 -4.37 -1.47 -1.82 -6.1 -4.3 -2.97 3.34 -2.05 2.7 1.91 2.48 2.82 1.5
2000 1 -2.93 -4.38 -3.84 -1.86 -0.54 -6.13 -4.71 -2.53 4.99 1.46 3.41 2.23 2.67 3.83 2.19
1.25 -2.71 -4.32 -3.81 -0.9 -0.84 -6.01 -4.83 -2.42 6.45 1.47 4.17 2.2 2.8 5.03 3.25
1.5 -2.59 -4.18 -3.6 0.58 2.09 -5.97 -4.71 -1.2 7.99 4.75 5.38 2.7 4.07 6.91 6.16
1.75 -2.63 -4.05 -3.17 0.5 2.4 -5.81 -4.59 -0.87 9.05 7.45 6.38 3.15 7.05 8.17 6.99
2 -2.55 -4.16 -3.68 -0.68 2.24 -5.93 -4.07 -0.42 9.46 7.39 6.66 3.16 4.94 8.15 6.7
0 -2.41 -4.18 -3.67 -0.93 -0.74 -5.66 -3.32 -3.41 0.01 0 0.76 0.76 -0.19 0.05 0.04
0.25 -1.97 -3.62 -2.37 -0.86 -0.19 -4.96 -3.18 -2.97 0.92 -1.51 1.49 1.12 0.5 0.85 0.25
0.5 -2.29 -3.76 -2.51 -0.9 -0.37 -5.05 -3.22 -2.6 1.96 -1.64 2.09 1.63 1.93 1.68 0.62
0.75 -2.22 -3.86 -2.73 -0.37 0.15 -5.2 -3.88 -2.56 3.4 -1.98 2.76 1.95 2.48 2.87 1.29
3500 1 -2.4 -4.04 -3.39 0.49 1.74 -5.5 -4.43 -2.25 5.18 2.6 3.47 2.14 2.38 4.09 3.05
1.25 -2.61 -4.21 -3.37 -0.77 0.26 -5.54 -4.78 -2.08 6.79 1.54 4.43 2.45 3.98 5.49 3.67
1.5 -2.23 -3.93 -2.59 1.03 1.9 -5.34 -4.41 -1.32 7.85 1.72 5.14 2.47 4.67 6.2 3.96
1.75 -2.46 -4.1 -3.17 0.71 2.59 -5.43 -4.63 -1.11 8.82 3.93 5.92 2.82 5.85 7.33 5.58
2 -2.64 -4.41 -3.73 -0.5 2.05 -5.81 -4.86 -1.17 9.21 6.79 6.58 2.73 6.2 8.27 6.16
0 -2.12 -3.78 -2.52 -0.64 -0.02 -5 -2.91 -2.89 -0.05 -0.04 0.74 0.74 0.01 -0.04 -0.06
0.25 -2.25 -3.9 -3.11 -0.24 -0.55 -5.14 -2.91 -2.93 0.81 -1.68 1.49 1.04 0.97 0.73 0.12
0.5 -2.06 -3.69 -2.5 0.1 -0.16 -5 -3.3 -2.53 1.88 -1.81 2.04 1.47 1.42 1.67 0.59
0.75 -2.3 -4.07 -3 -0.74 0.02 -5.25 -3.8 -2.57 3.42 -1.18 2.73 2.03 2.2 2.78 1.22
5000 1 -2.22 -3.95 -2.86 -0.2 0.41 -5.23 -4.33 -2.57 4.96 1.25 3.51 2.12 3.17 4.12 2.68
1.25 -2.52 -4.16 -3.41 0.4 0.95 -5.46 -4.72 -1.87 6.89 0.74 4.52 2.43 4.1 5.65 4.13
1.5 -2.33 -3.97 -2.96 0.17 1.78 -5.1 -4.55 -1.32 8.15 1.48 5.42 2.58 5.85 6.88 4.61
1.75 -1.86 -3.47 -2.45 1.03 2.74 -4.64 -4.07 -0.39 8.89 6.29 6.15 2.98 6.81 7.63 6.15
2 -2.25 -3.9 -3.03 1.32 3.8 -5.16 -4.27 -0.57 9.51 6.84 6.61 3.26 7.45 8.2 6.85

### v2

@@ Results
Train Shift KL CForest S-learner DR-learner T-learner
size degree Div. OS wOS OS wOS OS Ridge Tree OS wOS Ridge Tree
0 0.07 -0.59 -0.66 -2.25 -2.27 -2.28 -2.32 -2.42 -4.59 -4.6 -3.73 -4.4
0.25 0.22 -2.94 -1.19 -4.51 -3.91 -3.96 -3.38 -3.88 -6.21 -5.58 -4.39 -5.64
0.5 0.60 -5.21 -1.94 -6.68 -5.78 -5.14 -5.16 -5.36 -7.94 -6.79 -6.71 -8.1
0.75 1.05 -7.23 -2.63 -8.54 -7.44 -6.94 -5.77 -6.43 -9.59 -7.92 -8.42 -9.96
500 1 1.47 -8.57 -4.95 -10.07 -8.73 -8.58 -6.96 -7.66 -11.09 -9.13 -9.26 -10.79
1.25 1.76 -9.07 -5.13 -10.71 -9.44 -6.83 -6.66 -7.65 -11.62 -9.41 -10.17 -11.92
1.5 2.02 -9.56 -6.54 -11.14 -9.84 -8.59 -6.72 -8.72 -12 -9.49 -10.49 -12.25
1.75 2.21 -9.71 -5.74 -11.54 -10.19 -8.4 -7.83 -8.57 -12.48 -9.76 -11.2 -12.58
2 2.35 -10.04 -7.48 -11.7 -10.16 -9.39 -7.1 -8.24 -12.63 -9.29 -10.75 -12.02
0 0.02 0.18 0.17 -2.05 -2.06 -1.62 -1.63 -1.48 -3.51 -3.52 -3.3 -3.46
0.25 0.16 -2.01 -0.22 -3.93 -3.37 -2.49 -2.87 -3.04 -4.83 -4.28 -4.3 -4.78
0.5 0.53 -3.88 -0.64 -5.86 -4.88 -3.36 -3.81 -3.81 -6.53 -5.42 -6.2 -7.06
0.75 0.99 -5.55 -1.76 -7.54 -6.37 -4.39 -3.9 -4.55 -8.12 -6.72 -7.29 -8.26
2000 1 1.39 -6.5 -1.8 -8.64 -7.27 -5.98 -5.41 -5.71 -9.15 -7.28 -8.18 -9.26
1.25 1.73 -7.18 -2.12 -9.45 -8.04 -6.11 -5.29 -5.73 -9.95 -7.68 -8.98 -10.53
1.5 1.93 -7.6 -2.23 -9.74 -8.27 -6.07 -6.02 -6.41 -10.26 -7.87 -9.3 -10.63
1.75 2.09 -7.8 -3.98 -10.05 -8.81 -6.2 -5.67 -6.62 -10.55 -8.44 -9.95 -11.77
2 2.24 -7.87 -2.7 -10.19 -8.6 -6.84 -5.42 -6.29 -10.63 -7.7 -9.79 -11.07
0 0.01 0.13 0.16 -2.01 -2 -1.41 -1.46 -1.54 -3.17 -3.17 -3.02 -3.15
0.25 0.15 -1.64 -0.01 -3.64 -3.09 -2.29 -2.4 -2.35 -4.37 -3.86 -4.1 -4.39
0.5 0.52 -3.5 -0.4 -5.53 -4.53 -3.6 -3.47 -3.77 -6.08 -5.04 -5.9 -6.55
0.75 0.98 -4.77 -0.37 -6.87 -5.61 -4.15 -4.01 -4.16 -7.27 -5.73 -6.72 -7.8
3500 1 1.38 -5.83 -0.64 -8.12 -6.73 -5.24 -4.64 -4.55 -8.57 -6.61 -7.76 -8.58
1.25 1.70 -6.39 -0.74 -8.74 -7.25 -4.73 -5.2 -6.17 -9.08 -7.02 -8.4 -9.62
1.5 1.92 -6.7 -0.03 -9.06 -7.42 -6.11 -5.35 -5.37 -9.5 -7.08 -8.98 -10
1.75 2.10 -7 -0.6 -9.29 -7.63 -6.05 -5.85 -5.81 -9.68 -7.09 -8.8 -10.08
2 2.24 -7.17 -1.68 -9.58 -7.77 -6.13 -6.14 -6.2 -10.07 -7.46 -9.1 -9.93
0 0.01 0.21 0.22 -1.91 -1.89 -1.2 -1.35 -1.23 -3 -2.99 -2.96 -3.28
0.25 0.15 -1.47 0.09 -3.49 -2.94 -2.01 -2.14 -2.18 -4.14 -3.67 -3.93 -4.63
0.5 0.52 -3.05 0.05 -5.08 -4.04 -2.9 -2.94 -3.08 -5.53 -4.47 -5.01 -5.58
0.75 0.97 -4.53 -0.21 -6.68 -5.4 -3.9 -3.92 -4 -7.11 -5.58 -6.81 -7.51
5000 1 1.40 -5.38 0.18 -7.65 -6.25 -4.79 -4.55 -4.7 -8.02 -6.14 -7.51 -8.49
1.25 1.69 -5.99 -0.97 -8.28 -6.76 -4.67 -5.11 -4.95 -8.64 -6.45 -8.24 -9.2
1.5 1.93 -6.32 -0.21 -8.76 -6.99 -5.27 -4.71 -5.39 -9.11 -6.76 -8.62 -9.71
1.75 2.08 -6.42 -0.64 -8.83 -7.1 -5.64 -4.39 -5.24 -9.28 -6.77 -8.57 -9.49
2 2.23 -6.88 -2.61 -9.24 -7.47 -6.14 -5.33 -5.35 -9.68 -7.28 -9.12 -10.24
@@
Train Shift KL CForest S-learner DR-learner T-learner
size degree Div. OS wOS OS wOS OS Ridge Tree OS wOS Ridge Tree
0 0.07 48.96 48.83 61.44 61.47 49.01 36.24 41.68 50.37 50.47 37.92 48.71
0.25 0.22 58.03 52.83 79.93 79.13 64.22 47.74 53.11 72.51 64.08 52.5 67.26
0.5 0.60 75.47 59.45 107.39 107.74 76.5 65.54 69.39 103.71 86.43 80.42 111.06
0.75 1.05 98.37 78.02 134.31 132.05 109.67 83.42 87.06 132.66 105.15 107.44 147.33
500 1 1.47 116.47 94.48 160.77 154.28 126.34 110.72 111.87 164.12 127.02 121.08 165.23
1.25 1.76 121.81 132.84 169.13 165.05 142.91 108.59 128.54 173.35 133.51 138.64 186.45
1.5 2.02 128.21 144.18 177.03 169.83 133.75 115.82 119.92 179.79 137.04 142.15 194.81
1.75 2.21 129.24 139.23 183.05 177.42 153.78 109.25 123.72 189.41 140.58 158.14 198.73
2 2.35 136.23 156.07 187.09 176.18 143.81 114.75 133.21 193.61 136.73 148.47 192.43
0 0.02 31.25 31.29 41.25 41.46 26.21 23.87 25.76 32.69 32.76 28.04 33.36
0.25 0.16 35.78 33.1 56.21 54.81 37.36 31.1 35.15 49.45 42.36 40.36 51.18
0.5 0.53 47.93 36.68 78.99 74.7 53.38 47.29 58.23 74.61 57.9 68 86.61
0.75 0.99 62.81 43.77 102.92 98.62 67.95 55.98 59.31 101.23 78.57 83.52 107.64
2000 1 1.39 71.52 50.4 118.4 112.52 76.93 65.76 73.73 117.39 86.09 96.77 125.88
1.25 1.73 78.65 65.43 131.65 125.92 75.88 68 82.42 130.8 91.11 111.73 148.89
1.5 1.93 83.4 87.66 134.23 128.7 80.8 77.14 94 134.91 95.45 113.56 149.28
1.75 2.09 84.4 80.99 139.06 138.6 86.48 81.35 96.41 139.34 104.81 125.67 171.48
2 2.24 85.39 117.33 139.58 135.01 85.56 82.29 84.27 139.53 90.69 120.89 155.25
0 0.01 26.41 26.48 34.62 34.77 22.3 20.45 22.49 27.94 27.94 24.09 28.1
0.25 0.15 29.68 27.89 48.19 46.15 29.57 27.95 28.28 42.62 36.47 36.65 43.35
0.5 0.52 40.64 32.07 69.51 64.46 40.27 44.18 42.4 66.82 52.49 62.16 76.77
0.75 0.98 50.62 34.89 86.45 80.05 55.44 50.07 52.88 83.88 60.45 72.47 96.5
3500 1 1.38 60.8 40.59 105.13 99.35 58.89 55.2 65.03 104.92 73.67 88.04 109.7
1.25 1.70 65.42 53.8 113.57 105.53 73.34 60.93 71.11 112.44 79.46 98.13 128.19
1.5 1.92 67.51 58.93 115.71 105.15 66.37 63.82 62.2 117.23 78.11 106 131.51
1.75 2.10 71.14 84.67 119.63 110.99 68.86 60.95 72.15 119.74 79.36 100.75 133.81
2 2.24 72.57 94.48 124.32 112.8 70.37 68.83 68.96 127.26 84.85 106.19 128.92
0 0.01 23.76 23.84 30.99 31.12 19.81 18.33 19.01 25.41 25.37 22.87 28.73
0.25 0.15 26.53 25.12 43.11 40.92 25.98 25.28 26.89 38.97 33.2 36.33 45.83
0.5 0.52 35.34 28.45 60.29 54.75 32.56 31.91 35.93 57.47 42.99 48.56 60.4
0.75 0.97 46.53 31.58 81.14 73.84 48.41 41.51 53.5 81.54 58.2 75.15 91.08
5000 1 1.40 53.35 35.06 94.19 86.83 53.36 52.64 59.22 94.05 64.56 83.78 105.81
1.25 1.69 58.87 41.09 102.34 93.53 54.68 54.36 56.8 102.96 68.77 94.5 117.64
1.5 1.93 61.59 51.86 109.6 96.23 57.74 59.66 65.16 110.6 73.29 99.92 125.89
1.75 2.08 61.85 64.77 108.83 97.83 65.31 62.42 68.62 111.54 71.68 97.3 120.56
2 2.23 67.63 65.96 116.97 105 67.87 60.4 78.9 119.41 81.65 107.55 134.97
@@
Train Shift KL CForest S-learner DR-learner T-learner CForest S-learner DR-l T-learner
size degree Div. OS OS OS Ridge Tree OS Ridge Tree RCT wRCT RCT wRCT RCT RCT wRCT
0 2.82 -3.62 -4.41 -5.87 -3.91 -4.45 -7.43 -4.06 -4.23 0.01 0 0.83 0.82 -0.07 0.05 0.05
0.25 2.94 -3.75 -4.43 -5.1 -4.29 -3.88 -7.52 -4.18 -4.16 0.85 -0.64 1.54 1.25 0.64 0.89 0.41
0.5 3.30 -3.59 -4.39 -4.91 -4.73 -4.16 -7.32 -4.36 -3.89 1.87 -1.05 2.08 1.57 1.75 1.67 0.84
0.75 4.07 -3.41 -3.99 -4.4 -2.55 -3.6 -6.73 -4.24 -3.32 3.2 -1.37 2.7 1.88 2.56 2.67 1.17
500 1 5.44 -3.61 -4.38 -6.05 -2.71 -1.57 -7.53 -5.08 -3.25 5.01 1.15 3.5 2.13 1.71 4.13 2.17
1.25 6.95 -3.44 -4.12 -4.73 -3.54 -3.64 -7.13 -4.95 -2.56 6.56 0.48 4.43 2.39 3.17 5.47 3.83
1.5 9.37 -3.49 -4.16 -4.96 -3.61 -3.02 -7.31 -4.79 -1.77 7.89 2.07 5.16 2.53 4.42 6.26 3.97
1.75 12.16 -3.45 -4.18 -6.22 -3.01 -3.7 -6.98 -5.08 -1.35 8.81 5.85 6.01 2.65 5.95 7.47 5.69
2 15.04 -3.58 -4.36 -4.98 -3.1 -3.43 -7.42 -4.7 -0.95 9.5 2.38 6.6 2.83 7.1 8.21 6.07
0 2.52 -2.73 -4.23 -3.45 -1.89 -1.27 -5.95 -3.73 -3.61 0.01 0.12 0.78 0.81 0.08 0.03 0.06
0.25 2.92 -2.59 -4.07 -3.36 -1.23 -1.4 -5.78 -3.19 -3.17 0.79 -1.29 1.44 1.08 0.46 0.72 0.16
0.5 3.38 -2.73 -4.39 -3.53 -1.59 -1.65 -6.07 -3.76 -3.24 1.8 -2.16 2.07 1.57 1.25 1.55 0.6
0.75 4.09 -2.81 -4.4 -4.37 -1.47 -1.82 -6.1 -4.3 -2.97 3.34 -2.05 2.7 1.91 2.48 2.82 1.5
2000 1 5.26 -2.93 -4.38 -3.84 -1.86 -0.54 -6.13 -4.71 -2.53 4.99 1.46 3.41 2.23 2.67 3.83 2.19
1.25 6.98 -2.71 -4.32 -3.81 -0.9 -0.84 -6.01 -4.83 -2.42 6.45 1.47 4.17 2.2 2.8 5.03 3.25
1.5 9.43 -2.59 -4.18 -3.6 0.58 2.09 -5.97 -4.71 -1.2 7.99 4.75 5.38 2.7 4.07 6.91 6.16
1.75 11.93 -2.63 -4.05 -3.17 0.5 2.4 -5.81 -4.59 -0.87 9.05 7.45 6.38 3.15 7.05 8.17 6.99
2 15.22 -2.55 -4.16 -3.68 -0.68 2.24 -5.93 -4.07 -0.42 9.46 7.39 6.66 3.16 4.94 8.15 6.7
0 2.93 -2.41 -4.18 -3.67 -0.93 -0.74 -5.66 -3.32 -3.41 0.01 0 0.76 0.76 -0.19 0.05 0.04
0.25 2.85 -1.97 -3.62 -2.37 -0.86 -0.19 -4.96 -3.18 -2.97 0.92 -1.51 1.49 1.12 0.5 0.85 0.25
0.5 3.39 -2.29 -3.76 -2.51 -0.9 -0.37 -5.05 -3.22 -2.6 1.96 -1.64 2.09 1.63 1.93 1.68 0.62
0.75 4.01 -2.22 -3.86 -2.73 -0.37 0.15 -5.2 -3.88 -2.56 3.4 -1.98 2.76 1.95 2.48 2.87 1.29
3500 1 5.27 -2.4 -4.04 -3.39 0.49 1.74 -5.5 -4.43 -2.25 5.18 2.6 3.47 2.14 2.38 4.09 3.05
1.25 6.79 -2.61 -4.21 -3.37 -0.77 0.26 -5.54 -4.78 -2.08 6.79 1.54 4.43 2.45 3.98 5.49 3.67
1.5 9.18 -2.23 -3.93 -2.59 1.03 1.9 -5.34 -4.41 -1.32 7.85 1.72 5.14 2.47 4.67 6.2 3.96
1.75 11.94 -2.46 -4.1 -3.17 0.71 2.59 -5.43 -4.63 -1.11 8.82 3.93 5.92 2.82 5.85 7.33 5.58
2 15.15 -2.64 -4.41 -3.73 -0.5 2.05 -5.81 -4.86 -1.17 9.21 6.79 6.58 2.73 6.2 8.27 6.16
0 2.64 -2.12 -3.78 -2.52 -0.64 -0.02 -5 -2.91 -2.89 -0.05 -0.04 0.74 0.74 0.01 -0.04 -0.06
0.25 2.92 -2.25 -3.9 -3.11 -0.24 -0.55 -5.14 -2.91 -2.93 0.81 -1.68 1.49 1.04 0.97 0.73 0.12
0.5 3.33 -2.06 -3.69 -2.5 0.1 -0.16 -5 -3.3 -2.53 1.88 -1.81 2.04 1.47 1.42 1.67 0.59
0.75 4.18 -2.3 -4.07 -3 -0.74 0.02 -5.25 -3.8 -2.57 3.42 -1.18 2.73 2.03 2.2 2.78 1.22
5000 1 5.23 -2.22 -3.95 -2.86 -0.2 0.41 -5.23 -4.33 -2.57 4.96 1.25 3.51 2.12 3.17 4.12 2.68
1.25 6.87 -2.52 -4.16 -3.41 0.4 0.95 -5.46 -4.72 -1.87 6.89 0.74 4.52 2.43 4.1 5.65 4.13
1.5 9.18 -2.33 -3.97 -2.96 0.17 1.78 -5.1 -4.55 -1.32 8.15 1.48 5.42 2.58 5.85 6.88 4.61
1.75 12.10 -1.86 -3.47 -2.45 1.03 2.74 -4.64 -4.07 -0.39 8.89 6.29 6.15 2.98 6.81 7.63 6.15
2 15.23 -2.25 -3.9 -3.03 1.32 3.8 -5.16 -4.27 -0.57 9.51 6.84 6.61 3.26 7.45 8.2 6.85